OpenStudy (pottersheep):

Max/min question grade 10...just cheking answers

OpenStudy (pottersheep):

A lifeguard marks off a rectangular swimming area at a beach with 200m of rope. What is the greatest area of water she can enclose?

OpenStudy (pottersheep):

I got 2,500....but the answer sheet says 5,000! Which one is right?

OpenStudy (pottersheep):

Please?

OpenStudy (turingtest):

Perimiter=P P=2(l+w) wait, do we have to use calculus?

OpenStudy (turingtest):

Perimiter=P P=2(L+w)=200 100=L+w L=100-w Area is maximized when the derivative is zero: A=wL=w(100-w)=100w-w^2 A'=0=100-2w w=50 from earlier we have L=100-w=100-50=50 so the greatest possible area is A=wL=(50)(50)=2500

OpenStudy (turingtest):

Yep your sheet is wrong, I'd bet on it.

OpenStudy (pottersheep):

Thanks so much